OpenStudy (anonymous):

A box contains 100 colored chips; some are yellow and some are green. John chooses a chip at random, records the color, and places it back in the bag. John has recorded 57 yellow chips and 19 green chips. Using these results, what is the predicted number of green chips in the box? @sammsn8801

OpenStudy (anonymous):

@samman8801

OpenStudy (anonymous):

@RebeccaPink6398

OpenStudy (anonymous):

@itsbribro

OpenStudy (anonymous):

WILL FAN AND MEDAL

OpenStudy (anonymous):

so 43? @samman8801

OpenStudy (anonymous):

the answer choices are 25 43 76 81

OpenStudy (anonymous):

@samman8801

OpenStudy (anonymous):

@RebeccaPink6398

OpenStudy (anonymous):

I think @amistre64 Would be a much bigger help then me, He helped me earlier, im sure he will help you

OpenStudy (anonymous):

@amistre64 please help

OpenStudy (anonymous):

Because im not even positive If I did it right and I dont want you to get the answer wrong

OpenStudy (anonymous):

tell me what you think it is and thats nice that you dont want me to get it wrong!

OpenStudy (anonymous):

@samman8801

OpenStudy (amistre64):

what would we expect from the ratio we have already?

OpenStudy (anonymous):

im not sure @amistre64 i suck at math lol

OpenStudy (amistre64):

im not sure isnt a valid answer, it comes across as lazy ... what have been your attempts?

OpenStudy (anonymous):

i think its 43 but i dont know and i wanna get it right @amistre64

OpenStudy (amistre64):

well, tell me why you think its 43, what process led you there?

OpenStudy (anonymous):

What I wanted to do was add the numbers you had and subtract it by 100 but I got 24 and Ya I wasnt sure either

OpenStudy (anonymous):

i subtracted 57 from 100 and got 43 @amistre64

OpenStudy (amistre64):

adding is close, its a good direction, but a ratio is scaled 57 to 19 is our expected ratio 57k + 19k = 100 is our expected total of chips.

OpenStudy (anonymous):

the answer choices are 43 25 76 81 @amistre64

OpenStudy (amistre64):

lets say we actually have 10 yellow and 30 green our ratio is 1 to 3 1k + 3k = 40, what is k?

OpenStudy (anonymous):

10?

OpenStudy (amistre64):

right, so our process is valid 57k + 19k = 100 what is k?

OpenStudy (anonymous):

well i dont really understand because lets say k was 10 and you multiply 57 by 10 you pass 100 by a lot

OpenStudy (anonymous):

maybe k could be 2? @amistre64

OpenStudy (amistre64):

dont mix the example with the actual problem the example shows a process that can be validated. work the same process

OpenStudy (anonymous):

so maybe k could be 2 or 1?

OpenStudy (amistre64):

57k + 19k = 100 (57+19)k = 100 k = 100/76 so 19k is the number of green we would expect in the bag from the given ratio 19*100/76 is?

OpenStudy (anonymous):

43 or 24. I got them both @amistre64

OpenStudy (amistre64):

you cant get both of them from 19*100/76, its only one value.

OpenStudy (anonymous):

i got 43! A

OpenStudy (amistre64):

well, i dont agree with it, but good luck with it

OpenStudy (anonymous):

can we move onto another one? i got the answer for this next one but since your like super smart can you correct it?

OpenStudy (anonymous):

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 9 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section? 15 over 20 11 over 20 6 over 20 5 over 20

OpenStudy (anonymous):

i got d @amistre64

OpenStudy (amistre64):

how did you come to d? this is a study site, so method is important.

OpenStudy (anonymous):

ok so i looked at the fractions and i had done an exact problem like it before and 5 over 20 was the answer. but now that i look at it again i think it could be another one @amistre64

OpenStudy (amistre64):

red + blue + green = some amount we arent given the actual amount, but its save to say we can compare it to 20 9 of which are red, 6 of which are blue 9 + 6 + green = 20 green = 20 - 9 - 6

OpenStudy (amistre64):

we would expect 5 out of 20 is fine

OpenStudy (anonymous):

thank you! mind helping with more? @amistre64 you are a life saver

OpenStudy (amistre64):

in a new post, sure. this one is lagging my computer due to its length